Determine whether the lines l1 and l2 are parallel, perpendicular, or neither.
Line l1 goes through (3, –8) and (–3, 9). Line l2 goes through (–1, –5) and (–7, 12).

Answer by Alan3354(69443)

Determine whether the lines l1 and l2 are parallel, perpendicular, or neither.
Line l1 goes through (3, –8) and (–3, 9). Line l2 goes through (–1, –5) and (–7, 12).
--------
Find the slopes of the 2 lines. If the slopes are the same, they're parallel. If the slope of one (m) and the other is -1/m, they're perpendicular.
-----------
Line 1 thru (-3,9) and (3,-8). Always work "left to right", increasing x.
m1 = (y2-y1)/(x2-x1)
m1 = (-8-9)/(3 - (-3))
m1 = -17/6
-----------
Line 2 thru (-7,12) and (-1,-5)
m2 = (-5 -12)/(-1 -(-7))
m2 = -17/6
m2 = m1, so they're parallel
------
Working with increasing values of x is a convention, meaning something that's agreed on. Without that, the values of m could be inverted depending on the order a person chose.